OpenStudy (anonymous):

Is the diff equation linear? Separable? Neither? Both? (t^2 +1) dy/dt = yt -y

OpenStudy (anonymous):

(t^2 +1) dy/dt = yt -y=y(t-1)

OpenStudy (anonymous):

I see. Thanks